figures/purdue

A systems approach to dissect the tissue-specific architecture of cellular

networks

S. Mohammadi, B. Saberidokht, S. Subramaniam, \underline{A. Grama}

Department of Computer Science, Purdue University Department of Bioengineering, University of California, San Diego

Do I look like a microbe? figures/purdue

Outline

Do I look like a microbe?